At 497.8 °C, the Ke for the reaction: 2A + 5B # 4AB is 6.25 · 10° What is Keg? Enter your answer in scientific notation.
Answer #1

Keq = 1.58 × 106

2A 58 - 418 K8= 6.25 At (497. 8 + 273K Keg - (A874 - 770.8K [A] [875 Kp = (PAB) (PA) ² (18) As PV = nRT or p = nRT (P = CRT]